LOCAL MODE EFFECTS ON VIBRATIONAL AND ROTATIONAL ENERGY LEVEL PATTERNS IN SMALL SYMMETRICAL MOLECULES

\underline{L. HALONEN}, Laboratory of Physical Chemistry, P. O. Box 55 (A. I. Virtasen aukio 1), FIN-00014 University of Helsinki, Finland.